Advertisements
Advertisements
प्रश्न
A lottery with 600 tickets gives one prize of ₹ 200, four prizes of ₹ 100, and six prizes of ₹ 50. If the ticket costs is ₹ 2, find the expected winning amount of a ticket
Advertisements
उत्तर
Given, total number of tickets = 600
Prizes to be given: One prize of Rs. 200
Four prizes of ₹ 100
Six prizes of ₹ 50
Let ‘X’ be the random variable denotes the winning amount and it can take the values 200, 100 and 50.
Probability of winning ₹ 200 = `1/600`
Probability of winning ₹ 100 = `4/600`
Probability of winning ₹ 50 = `6/600`
∴ Probability mass function is
| x | 200 | 100 | 50 |
| F(x) | `1/600` | `4/600` | `6/600` |
∴ E(X) = 200 `sum x f(x)`
= `200/600 + 400/600 + 300/600`
= `900/600`
= 1.5
Expected winning amount = Amount won – Cost of lottery
= 1.50 – 2.00
= – 0.50
i.e., Loss of ₹ 0.50
APPEARS IN
संबंधित प्रश्न
For the random variable X with the given probability mass function as below, find the mean and variance.
`f(x) = {{:(1/10, x = 2"," 5),(1/5, x = 0"," 1"," 3"," 4):}`
For the random variable X with the given probability mass function as below, find the mean and variance.
`f(x) = {{:((4 - x)/6, x = 1"," 2"," 3),(0, "otherwise"):}`
Choose the correct alternative:
Four buses carrying 160 students from the same school arrive at a football stadium. The buses carry, respectively, 42, 36, 34, and 48 students. One of the students is randomly selected. Let X denote the number of students that were on the bus carrying the randomly selected student. One of the 4 bus drivers is also randomly selected. Let Y denote the number of students on that bus. Then E(X) and E(Y) respectively are
Choose the correct alternative:
On a multiple-choice exam with 3 possible destructive for each of the 5 questions, the probability that a student will get 4 or more correct answers just by guessing is
Choose the correct alternative:
If P(X = 0) = 1 – P(X = 1). If E[X] = 3 Var(X), then P(X = 0) is
What are the properties of Mathematical expectation?
How do you defi ne variance in terms of Mathematical expectation?
A person tosses a coin and is to receive ₹ 4 for a head and is to pay ₹ 2 for a tail. Find the expectation and variance of his gains
Choose the correct alternative:
Value which is obtained by multiplying possible values of a random variable with a probability of occurrence and is equal to the weighted average is called
Choose the correct alternative:
Probability which explains x is equal to or less than a particular value is classified as
Choose the correct alternative:
If X is a discrete random variable and p(x) is the probability of X, then the expected value of this random variable is equal to
Choose the correct alternative:
Which of the following is not possible in probability distribution?
Choose the correct alternative:
A discrete probability distribution may be represented by
Choose the correct alternative:
E[X – E(X)] is equal to
Choose the correct alternative:
`int_(-oo)^oo` f(x) dx is always equal to
Prove that V(aX) = a2V(X)
Prove that V(X + b) = V(X)
The time to failure in thousands of hours of an important piece of electronic equipment used in a manufactured DVD player has the density function
f(x) = `{{:(2"e"^(-2x)",", x > 0),(0",", "otherwise"):}`
Find the expected life of this piece of equipment
